Methanol is manufactured industrially by the reaction CO(g) + 2H2(g) <-> CH3OH(g). A gaseous mixture at 500K is 0.020 M CH3OH; 0.10M CO; and 0.10M H2. Calculate the equilibrium constant for the reation if the temperature is 300K. At what side will the equilibrium lie?

1
Expert's answer
2022-05-23T17:01:02-0400

CO(g) + 2 H2(g) <=> CH3OH(g).

Since the deltaH is positive, it means that the products have more energy than the reactants. This means that the reaction is endothermic.

in endothermic reactions, the chemical equilibrium shifts to the left as the temperature decreases.

Ke = [CH3OH] / [CO] * [H2]2

Ke = 0,02 / 0,1* 0,12 = 0,02 / 0,001 = 20

Answer: Ke = 20